OpenStudy (anonymous):

a^2+6ax-40x^2 factor out

myininaya (myininaya):

A=a^2 B=6ax C=-40x^2 what are two factors of A*C that add up to be B A*C=-40a^2x^2=(10ax)(-4ax) B=10ax-4ax so a^2+10ax-4ax-40x^2 now factor by grouping

OpenStudy (anonymous):

what do you mean by grouping

myininaya (myininaya):

(a^2+10ax)+(-4ax-40x^2)

myininaya (myininaya):

what do the terms in the first group have in common? what do the terms in the second group have in common?

OpenStudy (anonymous):

nothing...??

myininaya (myininaya):

?

myininaya (myininaya):

so there isn't an a in common in the first group? there isn't a -4x in common in the second group?

OpenStudy (anonymous):

theres not 2 -4x's so there isnt any -4x in common and theres a^2 and a but those are different they arent the same

myininaya (myininaya):

both terms in the first group DO have the factor a in common both terms in the second group DO have the factor -4x in common so we have a(a+10x)-4x(a+10x)

OpenStudy (anonymous):

so thats the final answer??

myininaya (myininaya):

we haven't factored it yet factor means numbers being multiplied those two terms are being subtracted (or you could say added) so what does the first term have in common with the second term now?

OpenStudy (anonymous):

10x and a???

myininaya (myininaya):

you mean 10x+a or a+10x right?

OpenStudy (anonymous):

suree

myininaya (myininaya):

ok so factored that out and tell me what you get?

OpenStudy (anonymous):

a^2+10x^2??

myininaya (myininaya):

no try again think can you factor this: 3ax+9x they both have in 3x in common so we can factor 3x out 3x(a+3)

OpenStudy (anonymous):

i dotn know how to do any of these i dont know what any of this means i have a 26% in my algebra class and i have about 17 more problems to do and none of this is making sense

myininaya (myininaya):

ok are problem above we are trying to factor a(a+10x)-4x(a+10x) both terms have a+10x in common so factor it out (a+10x)(a-4x) a was lefted from the first term -4x was lefted from the second term

OpenStudy (anonymous):

?????

myininaya (myininaya):

?????

myininaya (myininaya):

lol

OpenStudy (anonymous):

i have no clue what your talking about iwioth all this stuff matching up and how things are left out and what not

OpenStudy (anonymous):

i have no clue what your talking about iwioth all this stuff matching up and how things are left out and what not

myininaya (myininaya):

it sounds like you need an in-person tutor this is as well as i can explain it online and probably in person maybe you should review some easier forms of factoring first

myininaya (myininaya):

lol i actually try to explain it to you and others are just giving you answers and you give them medals lol